# SmileyLlama: Modifying Large Language Models for Directed Chemical Space Exploration

### v3

@@ 3.2 Property Specification
In Table 2 we show the average percentage of valid, distinct SMILES strings generated for a complete panel of families of tasks. This benchmark is distinct from other conditional molecule generation benchmarks16, 32 in that we are testing SmileyLlama’s ability to robustly generate molecules with properties in value ranges rather than molecules with a property of a specific value. This is for the reason of what is of interest to medicinal chemistry; many of the "rules of thumb", such as counting the number of Lipinski violations or number of hydrogen bond donors and acceptors, are based on value ranges during chemical exploration. Additionally, LLMs tend to struggle with numbers which have many degrees of precision and must be split into several tokens59. Hence we did not represent this category in the prompt during training.
Table 2: Percentage of valid, distinct generated molecules over a more complete panel of tasks using SmileyLlama. We performed a panel of 380 tasks (see Supplementary Information for complete definition) to assess how well SmileyLlama follows directions in a prompt. For each task, we generate 1000 molecules from a prompt requesting some property, and score the result based on the proportion of molecules which are valid, distinct, and satisfy the properties requested in the prompt. Finally, we collect some tasks into families (such as those with different range values) and average their scores to produce the results shown in the table below. All benchmarks were at a temperature $T=$1.0 and a maximum of 128 new tokens to increase computational efficiency. We also note that running SFT for two epochs rather than one did not seem to greatly affect the performance on these benchmarks.
SFT DPO
Present in Trained Prompt T=0.7 T=1.0 T=0.7 T=1.0
$\leq$ k H-bond donors 96.6 % 94.4 % 99.2 % 98.1 %
$\leq$ k H-bond acceptors 96.3 % 90.7 % 98.3 % 98.4 %
$\leq$ k Molecular weight 90.1 % 84.3 % 97.8 % 98.0 %
$\leq$ k ClogP 89.2 % 85.4 % 96.8 % 97.7 %
Rotatable bonds in range 90.0 % 86.7 % 94.5 % 94.0 %
Fraction $sp^{3}$ in range 87.6 % 85.0 % 96.3 % 95.9 %
TPSA in range 95.9 % 91.1 % 98.8 % 98.4 %
No bad SMARTS 92.5 % 89.6 % 94.9 % 94.3 %
Absence of macrocycle 98.4 % 95.3 % 98.4 % 97.3 %
Absence of warhead-related SMARTS 96.4 % 94.2 % 97.3 % 95.4 %
Lipinski rule-of-five 89.0 % 81.7 % 98.7 % 97.7 %
Presence of warhead-related SMARTS 58.3 % 51.0 % 74.9 % 73.0 %
Presence of Enamine substructures 51.0 % 51.5 % 67.7 % 70.1 %
Presence of macrocycle 38.8 % 44.7 % 58.0 % 57.8 %
Absent in Trained Prompt T=0.7 T=1.0 T=0.7 T=1.0
Rule-of-three 77.5 % 62.0 % 84.8 % 94.5 %
Exactly k H-bond donors 21.4% 19.7% 30.7% 30.4%
Exactly k H-bond acceptors 19.9% 14.4% 27.0% 31.9%
@@
Overall SmileyLlama model does very well on tasks on which it was trained through the engineered prompt. Although some of the properties were present in training, such as the Lipinski rule-of-5 and the presence of a macrocycle, both were more uncommon in the data and hence performance in these categories are less good. Similarly while there were many warhead-related SMARTS and Enamine substructures that were specified in the SFT training procedure, some were more common than others and performance is thus more middling as also seen in Table 2. Additionally, some tasks did not occur at all in the SFT training data, and yet are still parsible as valid SMILES. Finally, and as expected, SmileyLlama does poorly on tasks involving exact numerical specifications. We note that one has a choice to use SmileyLlama using lower temperatures at inference that can improve the SFT predictions as also seen in Table 2. SmileyLlama also performs well in driving not just singular molecular design tasks, but on compound tasks such as generating molecules similar to existing leads, i.e. “scaffold hopping” R-group modification, and/or structure-based design to grow molecules from ligand fragments. Figure 3a is an example of SmileyLlama model’s ability to generate molecules from all 320 substructures in the Enamine database61 that follow the Lipinski rule-of-five62, which encompasses most of the molecular properties with ranges listed in Table 2.
Figure 3: Conditional generation with SmileyLlama for fragment growth and before and after DPO compared to ChEMBL. (a) Example molecules generated by growing from one of the Enamine substructures and to satisfy Lipinski’s Rule of 5 using the prompt Output a SMILES string for a drug like molecule with the following properties: a substructure of O=C(O)c1ccc(C(F)(F)F)cc1, <= 500 MW, <=5 logP, <= 5 H-bond donors, <= 10 H-bond acceptors. (b) distribution of four properties satisfying Lipinski’s rule of five comparing ChEMBL molecules (orange) with molecules generated by SmileyLlama (blue) with the prompt Output a SMILES string for a drug like molecule with the following properties: <= 5 H-bond donors, <= 10 H-bond acceptors, <= 500 MW, <= 5 logP, compared to 1000 molecules generated by SmileyLlama with the same prompt after DPO (gray). MW and LogP distributions were estimated using a gaussian kernel density estimator (KDE).60 All results generated 1000 molecules at a temperature $T=$1.0 and a maximum of 128 new tokens.
@@
While the SFT SmileyLlama model typically does well on tasks on which it was trained through the engineered prompt, and can still perform adequately when queried with prompts different from ones on which it was trained (Table 2), we can further optimize SmileyLlama for specific tasks using DPO. DPO’s most popular application has been in improving the responses of LLM-derived chatbots, such as the instruct-tuned Llama3 models, but it has also found use in improving the outputs of CLMs 63 and avoiding the need to separately train a reward model 39. Here the relevance of DPO provides a way to further optimize the model by pairing desirable responses with undesirable responses. The model’s weights are then updated to be more likely to produce the ‘winner’ of the pairing and less likely to produce the ‘loser’ of the pairing. We generated our dataset by simply pairing up unsuccessful attempts at generating structures with successful attempts randomly for each task in Table 2. We concatenate these pairings into a single dataset and optimize for 1 epoch using DPO. SmileyLlama optimized with DPO significantly improved adherence to the prompt across nearly all tasks as seen in Table 2 and Figure 3b. Note that while DPO does cause the model to more robustly obey the rules in the prompt, it also shifts and narrows the property distribution compared to the training set as seen in Figure 3b and appears to be largely insensitive to temperature. SmileyLlama without DPO, on the other hand, occasionally does not obey the prompt but more faithfully reproduces the distribution of properties found in a filtered ChEMBL that satisfy Lipinski’s rules. Hence, our investigation indicates that DPO cna be quite useful for some tasks, but can sometimes have unintended consequences.
@@
3.3 3D Optimization of Drugs in Protein Active Sites with SmileyLlama
@@
The tests performed in Sections 3.1 and 3.2 do not take advantage of the 3-dimensional structural information of a putative drug nor its shape and molecular compatibility with a target protein active site. Hence we use SmileyLlama to generate unique and valid ligands which undergo further optimization for binding to a specific protein target when embedded in the iMiner framework21. iMiner generates novel inhibitor molecules for target proteins by combining deep reinforcement learning64, 65 with real-time 3D molecular docking using AutoDock Vina66, thereby simultaneously creating chemical novelty while constraining molecules for shape and molecular compatibility with target active sites. More specifically, the iMiner generative model uses an Average Stochastic Gradient Descent Weigh-Dropped Long Short-Term Memory (AWD-LSTM)67 recurrent neural network that predicts probability of string tokens to concatenate on a molecular string representation until a complete molecule is generated. In the subsequent RL stage, 2000 molecules in each epoch (typically $\sim$ 50 epochs) are sent to AutoDock simulations in parallel, and the docking scores are circled back to the RL model to adjust its parameters so that molecules generated in the next iteration will have better scores while retaining drug-likeness.
@@
Given the ascendancy of attention mechanisms and transformer architectures, such as those inherent in SmileyLlama, we replaced the generative AWD-LSTM component of iMiner with SmileyLlama, and replaced the iMiner optimization algorithm, Proximal Policy Optimization (PPO)68 with DPO. To validate the effectiveness of SmileyLlama in the iMiner context, we generate inhibitor molecules for the popular SARS2 MPro as an example. We chose this system because of the readily available experimental 3D structures69, 70 which provides the necessary information needed for structure-based ligand design using SmileyLlama and comparisons to other successful inhibitor molecules.
@@
For the unconditional de novo generation case, SmileyLlama learns the user prompt Output a SMILES string for a drug like molecule with the following properties: High SARS2PRO which pertains to minimizing the AutoDock Vina score while maximizing the drug likeliness score ($S_{DL}$) of the original iMiner reward function21. Figure 4a compares the docking scores of the original iMiner algorithm against SmileyLlama as a function of epoch number and with number of generated molecules per iteration. We notice first an improved data efficiency compared to iMiner, in which SmileyLlama requires roughly one-quarter of the epochs to reach a similar level of improved docking score. Due to iMiner’s use of SELFIES as its molecular representation, its validity is unparalleled, however, in later epochs, diversity crashes which explains the sharpening peaks in later iterations (Figure 4) and quantified further in Supplementary Figure S2 against the Guacamol benchmarks. This simply reflects convergence in the docking score, i.e. there are fewer novel molecules as docking score reaches the highest values. By contrast, SmileyLlama maintains more diversity while greatly improving the docking score to iMiner (Figure 4a) with some degradation in validity compared to iMiner (Supplementary Figure S2) likely due to the use of the SMILES representation.
Figure 4b shows the property distributions of the final optimized set of novel molecules from SmileyLlama from the prompt Output a SMILES string for a drug like molecule with the following properties: High SARS2PRO. While the property distributions are satisfactory for parts of the Lipinski Rule of 5, such as number of hydrogen bond donors and
@@
Figure 4: Comparison of the shift in docking score distributions for iMiner compared to SmileyLlama over optimization epochs as illustrated for SARS2-MPro. (a) For iMiner, in later epochs diversity crashes which explains the sharpening peaks in later iterations. SmileyLlama with DPO (SL+DPO) enforces diversity throughout the optimizations (Algorithm S3), which accounts for the broad peaks, and shows superior data efficiency relative to iMiner. (b) We compare two different user prompts: Sars2Pro and Sars2Pro+Ro5. All results were generated with 1024 SMILES at a temperature of $T=$1.0 and a maximum of 128 new tokens.
@@
acceptors, the MW and logP results are not conforming to drug-like values. This indicates some inadequacy of the iMiner reward function, i.e. a non-optimal weighting of docking scores compared to $S_{DL}$. While a standard CLM such as iMiner would require a reweighting and/or new terms in the loss/reward function, other hyperparameter tuning, and expensive retraining, a unique advantage of SmileyLlama is that the distribution of generated molecules’ properties can be shifted using nothing more than prompt engineering, with no retraining required. Figure 4b shows that combining prompts such as Output a SMILES string for a drug like molecule with the following properties: High SARS2PRO, <= 5 H-bond donors, <= 10 H-bond acceptors, <= 500 molecular weight, <= 5 logP , which we will refer to as the High SARS2Pro+Ro5 prompt improves properties such as MW and logP and drug-likeness scores significantly with some expected loss in high docking scores since smaller molecules make fewer intermolecular interactions.
@@
Figure 5 provides a set of novel molecules from SmileyLlama docked in the MPro active site with the two engineered prompts High SARS2Pro and High SARS2Pro+Ro5. Two of the higher scoring molecules resemble the variations of the perampanel drug with the trefoil structure which are tested inhibitors optimized by the Jorgensen group.71 However, unlike the molecules from their study that consistently retained the central pyridinone ring71, SmileyLlama molecules have replaced the trefoil hub with the pyrimidine functional group (see Figure 5c). Higher docking scores are found for quite different drug scaffolds (Figure 5(a,b)), but in all cases there is no significant homology match found in the Therapeutic Target Database.72 This would indicate that the generative capabilities of SmileyLlama are robust and outside of the pre-trained Llama model. Finally, the proposed drugs are synthetically accessible73, as made clear from their on average SAsimilar-$\sim$ 3 as seen in Supplementary Table S2.

### v4

@@ 3.2 Property Specification
In Table 2 we show the average percentage of valid, distinct SMILES strings generated for a complete panel of families of tasks. This benchmark is distinct from other conditional molecule generation benchmarksmol-gen-gpt0, yu2024llasmol in that we are testing SmileyLlama’s ability to robustly generate molecules with properties in value ranges rather than molecules with a property of a specific value. This is for the reason of what is of interest to medicinal chemistry; many of the "rules of thumb", such as counting the number of Lipinski violations or number of hydrogen bond donors and acceptors, are based on value ranges during chemical exploration. Additionally, LLMs tend to struggle with numbers which have many degrees of precision and must be split into several tokensschwartz-etal-2024-numerologic. Hence we did not represent this category in the prompt during training.
We compare SmileyLlama, both before and after DPO, with a model resulting from an ablation study on the efficacy of prompting. We study this by removing all indications of molecular properties from all of the prompts in the dataset used to train SmileyLlama; each molecule from ChEMBL is a completion to the same prompt (the prompt used for SmileyLlama when no properties of a molecule are selected). We then run SFT with exactly the same hyperparameters as SmileyLlama. Finally, we perform the same analysis of how well this model follows properties given in its prompt as we ran with SmileyLlama both before and after DPO. We find that this model performs quite poorly in comparison to SmileyLlama in this benchmark, only achieving 90+% performance on three tasks. This becomes especially pronounced when the properties are rarely found in the data, such as the presence of a macrocycle or a warhead-related SMARTS pattern. The stark contrast in performances speaks to the necessity of our prompt engineering scheme; we cannot rely purely on the knowledge of the foundation model when fine-tuning for chemical tasks.
@@
Overall SmileyLlama model does very well on tasks on which it was trained through the engineered prompt, especially when contrasted with the model resulting from the "prompt ablation" experiment. Although some of the properties were present in training, such as the Lipinski rule-of-5 and the presence of a macrocycle, both were more uncommon in the data and hence performance in these categories are less good. Similarly while there were many warhead-related SMARTS and Enamine substructures that were specified in the SFT training procedure, some were more common than others and performance is thus more middling as also seen in Table 2. Additionally, some tasks did not occur at all in the SFT training data, and yet are still parsible as valid SMILES. Finally, and as expected, SmileyLlama does poorly on tasks involving exact numerical specifications. We note that one has a choice to use SmileyLlama using lower temperatures at inference that can improve the SFT predictions as also seen in Table 2.
Table 2: Percentage of valid, distinct generated molecules over a more complete panel of tasks using SmileyLlama. We performed a panel of 380 tasks (see Supplementary Information for complete definition) to assess how well SmileyLlama follows directions in a prompt. For each task, we generate 1000 molecules from a prompt requesting some property, and score the result based on the proportion of molecules which are valid, distinct, and satisfy the properties requested in the prompt. Finally, we collect some tasks into families (such as those with different range values) and average their scores to produce the results shown in the table below. All benchmarks were at a temperature T=T=1.0 and a maximum of 128 new tokens to increase computational efficiency. We also note that running SFT for two epochs rather than one did not seem to greatly affect the performance on these benchmarks.
SFT DPO Prompt Ablation
Present in Trained Prompt T=0.7 T=1.0 T=0.7 T=1.0 T=0.7 T=1.0
$\leq$ k H-bond donors 96.6% 94.4% 99.2% 98.1 % 93.8% 91.8%
$\leq$ k H-bond acceptors 96.3% 90.7% 98.3% 98.4 % 76.6% 65.4%
$\leq$ k Molecular weight 90.1% 84.3% 97.8% 98.0 % 62.4% 58.8%
$\leq$ k ClogP 89.2% 85.4% 96.8% 97.7 % 67.9% 64.9%
Rotatable bonds in range 90.0% 86.7% 94.5% 94.0 % 62.2% 58.5%
Fraction $sp^{3}$ in range 87.6% 85.0% 96.3% 95.9 % 30.7% 31.8%
TPSA in range 95.9% 91.1% 98.8% 98.4 % 89.2% 83.1%
No bad SMARTS 92.5% 89.6% 94.9% 94.3 % 87.3% 85.6%
Absence of macrocycle 98.4% 95.3% 98.4% 97.3 % 97.1% 94.8%
Absence of warhead-related SMARTS 96.4% 94.2% 97.3% 95.4 % 94.6% 92.9%
Lipinski rule-of-five 89.0% 81.7% 98.7% 97.7 % 73.6% 64.0%
Presence of warhead-related SMARTS 58.3% 51.0% 74.9% 73.0 % 0.2% 0.4%
Presence of Enamine substructures 51.0% 51.5% 67.7% 70.1 % 2.7% 1.9%
Presence of macrocycle 38.8% 44.7% 58.0% 57.8 % 2.1% 2.1%
Absent in Trained Prompt T=0.7 T=1.0 T=0.7 T=1.0 T=0.7 T=1.0
Rule-of-three 77.5% 62.0% 84.8% 94.5% 7.5% 5.5%
Exactly k H-bond donors 21.4% 19.7% 30.7% 30.4% 17.4% 16.8%
Exactly k H-bond acceptors 19.9% 14.4% 27.0% 31.9% 9.2% 8.8%
SmileyLlama also performs well in driving not just singular molecular design tasks, but on compound tasks such as generating molecules similar to existing leads, i.e. “scaffold hopping” R-group modification, and/or structure-based design to grow molecules from ligand fragments. Figure 3a is an example of SmileyLlama model’s ability to generate molecules from all 320 substructures in the Enamine databaseenamine_essential_libraries that follow the Lipinski rule-of-fiveLIPINSKI20013, which encompasses most of the molecular properties with ranges listed in Table 2.
@@
While the SFT SmileyLlama model typically does well on tasks on which it was trained through the engineered prompt, and can still perform adequately when queried with prompts different from ones on which it was trained (Table 2), we can further optimize SmileyLlama for specific tasks using DPO. DPO’s most popular application has been in improving the responses of LLM-derived chatbots, such as the instruct-tuned Llama3 models, but it has also found use in improving the outputs of CLMs park_preference_2023 and avoiding the need to separately train a reward model rafailov_direct_2024. Here the relevance of DPO provides a way to further optimize the model by pairing desirable responses with undesirable responses. The model’s weights are then updated to be more likely to produce the ‘winner’ of the pairing and less likely to produce the ‘loser’ of the pairing. We generated our dataset by simply pairing up unsuccessful attempts at generating structures with successful attempts randomly for each task in Table 2. We concatenate these pairings into a single dataset and optimize for 1 epoch using DPO.
SmileyLlama optimized with DPO significantly improved adherence to the prompt across nearly all tasks as seen in Table 2 and Figure 3b. Note that while DPO does cause the model to more robustly obey the rules in the prompt, it also shifts and narrows the property distribution compared to the training set as seen in Figure 3b and appears to be largely insensitive to temperature. SmileyLlama without DPO, on the other hand, occasionally does not obey the prompt but more faithfully reproduces the distribution of properties found in a filtered ChEMBL that satisfy Lipinski’s rules. In the context of drug discovery, SFT is primarily beneficial for early exploration of chemical space, whereas DPO is a type of constraint optimization that limits generated molecules to desired sub-classes specified by the user.
Figure 3: Conditional generation with SmileyLlama for fragment growth and before and after DPO compared to ChEMBL. (a) Example molecules generated by growing from one of the Enamine substructures and to satisfy Lipinski’s Rule of 5 using the prompt Output a SMILES string for a drug like molecule with the following properties: a substructure of O=C(O)c1ccc(C(F)(F)F)cc1, <= 500 MW, <=5 logP, <= 5 H-bond donors, <= 10 H-bond acceptors. (b) distribution of four properties satisfying Lipinski’s rule of five comparing ChEMBL molecules (orange) with molecules generated by SmileyLlama (blue) with the prompt Output a SMILES string for a drug like molecule with the following properties: <= 5 H-bond donors, <= 10 H-bond acceptors, <= 500 MW, <= 5 logP, compared to 1000 molecules generated by SmileyLlama with the same prompt after DPO (gray). MW and LogP distributions were estimated using a gaussian kernel density estimator (KDE).KernelDensityEstimators1992 All results generated 1000 molecules at a temperature T=T=1.0 and a maximum of 128 new tokens.
@@
3.3 Optimization of Binding Affinity to Protein Active Sites with SmileyLlama/iMiner
@@
The tests performed in Sections 3.1 and 3.2 do not take advantage of the 3-dimensional structural information of a putative drug nor its shape and molecular compatibility with a target protein active site. Hence we use SmileyLlama to generate unique and valid ligands which undergo further optimization for binding to a specific protein target when embedded in the iMiner frameworkiminer. iMiner generates novel inhibitor molecules for target proteins by combining deep reinforcement learningolivecrona2017molecular, popova2018deep with real-time 3D molecular docking using AutoDock VinaTrott2010, thereby simultaneously creating chemical novelty while constraining molecules for shape and molecular compatibility with target active sites. More specifically, the iMiner generative model uses an Average Stochastic Gradient Descent Weigh-Dropped Long Short-Term Memory (AWD-LSTM)awd-lstm recurrent neural network that predicts probability of string tokens to concatenate on a molecular string representation until a complete molecule is generated. In the subsequent RL stage, 2000 molecules in each epoch (typically $\sim$ 50 epochs) are sent to AutoDock simulations in parallel, and the docking scores are circled back to the RL model to adjust its parameters so that molecules generated in the next iteration will have better scores while retaining drug-likeness.
Given the ascendancy of attention mechanisms and transformer architectures, such as those inherent in SmileyLlama, we replaced the generative AWD-LSTM component of iMiner with SmileyLlama, and replaced the iMiner optimization algorithm, Proximal Policy Optimization (PPO)ppo-schulman with DPO. This is largely due to the high memory requirement of PPO; since DPO does not need to fit a reward model to the predicted and realized rewards it requires far less memory than PPO. This becomes even more true when tuning a large model compared to a small one, since generally the reward model trained is the same size as the language model used to generate strings.
@@
To validate the effectiveness of SmileyLlama in the iMiner context, we generate inhibitor molecules for the SARS-CoV-2 Main Protease(MPro) as an example. One route to development of antiviral treatments to Covid-19 is through inhibition of MPro, an enzyme whose function is essential to the SARS-CoV-2 lifecycle.jin2020structure We chose this system not only due to its current relevance, but also because of the readily available experimental 3D structuresjin2020structure, zhang2020crystal which provides the necessary information needed for structure-based ligand design using SmileyLlama and comparisons to other successful inhibitor molecules for this target.
@@
For the unconditional de novo generation case, SmileyLlama learns the user prompt Output a SMILES string for a drug like molecule with the following properties: High SARS2PRO which pertains to minimizing the AutoDock Vina score while maximizing the drug likeliness score ($S_{DL}$) of the original iMiner reward functioniminer. Figure 4a compares the docking scores of the original iMiner algorithm against SmileyLlama as a function of epoch number and with number of generated molecules per iteration. We notice first an improved data efficiency compared to iMiner, in which SmileyLlama requires roughly one-quarter of the epochs to reach a similar level of improved docking score. Due to iMiner’s use of SELFIES as its molecular representation, its validity is unparalleled, however, in later epochs, diversity crashes which explains the sharpening peaks in later iterations (Figure 4) and quantified further in Supplementary Figure S2 against the Guacamol benchmarks. This simply reflects convergence in the docking score, i.e. there are fewer novel molecules as docking score reaches the highest values. By contrast, SmileyLlama maintains more diversity while greatly improving the docking score to iMiner (Figure 4a) with some degradation in validity compared to iMiner. At 15 epochs, SmileyLlama has 93% validity, compared to iMiner’s 100% validity (Supplementary Figure S4) likely due to the fact that SmileyLlama represents molecules with SMILES strings while iMiner uses SELFIES.
@@
Figure 4: Comparison of the shift in docking score distributions for iMiner compared to SmileyLlama over optimization epochs as illustrated for SARS2-MPro. (a) For iMiner, in later epochs diversity crashes which explains the sharpening peaks in later iterations. SmileyLlama with DPO (SL+DPO) enforces diversity throughout the optimizations (Algorithm S3), which accounts for the broad peaks, and shows superior data efficiency relative to iMiner. (b) We compare two different user prompts: Sars2Pro and Sars2Pro+Ro5. All results were generated with 2000 valid SMILES at a temperature of T=T=1.0 and a maximum of 128 new tokens.
@@
Figure 4b shows the property distributions of the final optimized set of novel molecules from SmileyLlama from the prompt Output a SMILES string for a drug like molecule with the following properties: High SARS2PRO. While the property distributions are satisfactory for parts of the Lipinski Rule of 5, such as number of hydrogen bond donors and acceptors, the MW and logP results are not conforming to drug-like values. This indicates some inadequacy of the iMiner reward function, i.e. a non-optimal weighting of docking scores compared to $S_{DL}$. While a standard CLM such as iMiner would require a reweighting and/or new terms in the loss/reward function, other hyperparameter tuning, and expensive retraining, a unique advantage of SmileyLlama is that the distribution of generated molecules’ properties can be shifted using nothing more than prompt engineering, with no retraining required. Figure 4b shows that combining prompts such as Output a SMILES string for a drug like molecule with the following properties: High SARS2PRO, <= 5 H-bond donors, <= 10 H-bond acceptors, <= 500 molecular weight, <= 5 logP, which we will refer to as the High SARS2Pro+Ro5 prompt improves properties such as MW and logP and drug-likeness scores significantly with some expected loss in high docking scores since smaller molecules make fewer intermolecular interactions.
[…]